A train is 300 meters long through a tunnel that is 2900 meters long. The train travels at 600 meters per minute. How long is the time between the moment the front of the train enters the tunnel and the moment the train exits the tunnel?

I believe that it is 4 minutes and 55 milliseconds. Im not sure though, please get back.

Answer by Theo(13342)

the tunnel is 2900 meters long.
the train is 300 meters long.
in order to get completely through the tunnel so that the rear end of the train is out of the tunnel, the train has to travel 2900 + 300 = 3200 meters.
at 600 meters per minutes, it would take the train 3200 / 600 = 5 and 1/3 minutes.
1/3 of a minute = 20 seconds.
it would take the train 5 minutes and 20 seconds to clear the tunnel.